Find the terminal points P(x,y) on the unit circle determined by the given value t
1
Expert's answer
2021-08-16T14:08:57-0400

Given "t, -2\pi\leq t\leq2\pi." Then


"x=\cos t, y=\sin t"

 The terminal point "P(\cos t, \sin t)."





"t=\dfrac{3\pi}{4}, x=\cos(\dfrac{3\pi}{4})=-\dfrac{\sqrt{2}}{2}, y=\sin(\dfrac{3\pi}{4})=\dfrac{\sqrt{2}}{2}"

"P(-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})"



"t=-\dfrac{\pi}{3}, x=\cos(-\dfrac{\pi}{3})=\dfrac{1}{2}, y=\sin(-\dfrac{\pi}{3})=-\dfrac{\sqrt{3}}{2}"

"P(\dfrac{1}{2}, -\dfrac{\sqrt{3}}{2})"